OpenStudy (anonymous):

convert to celsius.use the formula c=5/9*(F+32). 204°c

OpenStudy (anonymous):

Do you not understand the formula?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

204=5/9*(F+32). solve

OpenStudy (anonymous):

Well the 204 is you c in the above formula ok So you plug that in \[\Large 240~celcius=\]

OpenStudy (anonymous):

\[\frac{ 5 }{ 9 }F+\frac{5}{9}(32)\]

OpenStudy (anonymous):

5F/9+160/9

OpenStudy (anonymous):

ok im dumb

OpenStudy (anonymous):

\[\Large 240=\frac{5}{9}F+\frac{5}{9}(32)\]

OpenStudy (anonymous):

400

OpenStudy (anonymous):

im right of not

OpenStudy (anonymous):

ok im dumb

OpenStudy (anonymous):

you got it!

OpenStudy (anonymous):

wait your question asked for 204 not 240 sorry

OpenStudy (anonymous):

95.6

OpenStudy (anonymous):

i got a red big x

OpenStudy (anonymous):

on it

OpenStudy (anonymous):

Are you sure? \[\Large cancell ~out~fraction~~~~~(9/5)204=(9/5)(5/9({F+32}))\]

OpenStudy (anonymous):

It was 204=5/9(F+32) before

OpenStudy (anonymous):

367.2=F+32 subtract 32 from both sides ?=F